OpenStudy (anonymous):

Help please im not sure im understanding this (x-4)(x+3) I have to simplify it

OpenStudy (anonymous):

Split the expression into four parts (FOIL) 1) x * x = x ^2 2) -4 * x = -4x 3) 3 * x = 3x 4) -4 * 3 = -12 Then add everything together: \[x ^{2} - 4x + 3x - 12\] \[x ^{2} - x - 12\]